# Auto-Compressing Networks

### v1

@@ 12 Conclusions
Our theoretical and empirical analyses demonstrate that ACNs alter gradient flow, imposing implicit layer-wise training dynamics and resulting in distinct representations compared to feedforward and residual architectures. In practice, this leads to 30–80% of upper layers becoming effectively redundant, enabling faster inference and reduced memory usage without sacrificing accuracy. Experiments across diverse modalities (vision, language) and architectures (ViTs, Mixers, BERT) further show that ACNs match or outperform residual baselines, while offering greater robustness to noise and low-data regimes, excelling in transfer learning, boosting pruning techniques and reducing catastrophic forgetting by up to 18%—all without specialized tuning, overall suggesting that they learn better representations despite using fewer parameters.
Concluding, Auto-Compressing Networks (ACNs), building on implicit regularization through architectural design insights, represent a promising step towards more self-adapting neural architectures that allocate resources based on the task at hand, while learning sparse yet robust representations. Future research could expand ACNs to self-supervised and multi-task settings, leveraging the pre-training and fine-tuning paradigm. ACNs also hold promise for generative tasks, reducing inference costs and energy consumption. Additionally, developing inference-time algorithms that dynamically adjust the number of layers per sample for optimal performance and efficiency is an intriguing direction for future work. Last but not least, ACNs are only one possible long-connection architecture out of the many that are worth investigating further.
13 Limitations
Due to resource constraints, our proposed architecture was evaluated solely on relatively small scale tasks; however, it demonstrated robust and promising performance across various modalities, datasets, and state-of-the-art architectures within this scope. To fully assess its potential and limitations, further testing on a broader range of tasks is essential. Additionally, applying our method in self-supervised and multi-task learning settings, such as training large-scale language models or multimodal models, represents a significant and exciting avenue for future research.
Another limitation is the increased training time observed with ACNs compared to traditional architectures with residual connections. While we partially addressed this issue by employing parameterizations similar to DiracNets, a more comprehensive solution to reduce training time remains an open question. Of course this could be both a blessing and a curse, as longer training times might lead to learning better representations. In any case, further research into training schedules and initialization schemes is needed to resolve this trade-off.
14 Broader Impact
Our work contributes to the development of more efficient and robust neural network architectures by drawing inspiration from biological processes. By enabling networks to identify and prune redundant layers, we aim to reduce computational, memory, and energy requirements during inference, which will have a significant impact with broader adoption of AI technology. Conceptually, this line of research could lead to network architectures with inherent System 1 and System 2 capabilities (Kahneman, 2011), where networks adaptively use fewer layers—analogous to fast thinking—for easier tasks, and engage more layers—resembling slow thinking—for more complex tasks.
However, as with any advancement in AI, there is a potential for misuse. More efficient models could be leveraged to deploy AI systems more broadly, including in areas with insufficient oversight or in applications that may infringe on privacy or other ethical considerations.
Appendix A Experimental Details
@@
CIFAR-10 - MLP Mixer: The MLP Mixers have 16 layers with a hidden size of 128. The patch size is 4 (the input is 32x32, 3 channels). The MLP dimension $D_{C}$ is $512$, while $D_{S}$ is $64$. We are using the AdamW optimizer Loshchilov (2017) with a maximum learning rate of $0.001$ and a Cosine Scheduler with Warmup. The batch size is 64.
@@
Continual Learning Experiments: We are using the same MLP-Mixer setup with the C-far-10 experiment (see above). We train for 10 epochs in each task, using AdamW with learning rate of 0.001 and a batch size of 64. For Synaptic Intelligence we use a coefficient $\lambda=1$.
Figure 11: Main concept: (a) Start from a neural network with $L$ layers either randomly initialized or pretrained. (b) Add residual long connections from each layer to the ouput of the network and sum them (also remove any existing short residual connections - if any). During training of the resulting ACN network the majority of the information (shown here as darker vs lighter circles) will naturally concentrate at the lower layers. (c) You may now safely remove the top (two in our example) layers during inference without any performance loss.
Appendix B Gradient Propagation equations derivation

### v2

@@ 8 Conclusion
Our theoretical and empirical analyses demonstrate that ACNs alter gradient flow, imposing implicit layer-wise training dynamics and resulting in distinct representations compared to feedforward and residual architectures. In practice, this leads to 30–80% of upper layers becoming effectively redundant, enabling faster inference and reduced memory usage without sacrificing accuracy. Experiments across diverse modalities (vision, language) and architectures (ViTs, Mixers, BERT) further show that ACNs match or outperform residual baselines, while offering greater robustness to noise and low-data regimes, excelling in transfer learning, and reducing catastrophic forgetting by up to 18%—all without specialized tuning, overall suggesting that they learn better representations despite using fewer parameters. A summary of the main results is shown in Table 6. Another practical advantage of the auto-compression property of ACNs is its potential to enhance or complement other compression and efficient inference techniques. Preliminary results demonstrating this synergy are presented in Appendices G and H.
Concluding, ACNs highlight the potential of architectural design as implicit regularization and pave the way toward self-adapting neural networks that allocate depth and capacity dynamically. Future work may extend ACNs to self-supervised, multi-task, and generative settings, as well as explore per-sample adaptive inference and other short vs long architectural variants. Some initial
9 Limitations and Broader Impact
Due to resource constraints, our evaluation of ACNs was limited to relatively small-scale tasks, though the architecture consistently performed well across modalities, datasets, and state-of-the-art baselines. Broader validation—including large-scale, self-supervised, and multi-task settings (e.g., language or multimodal models)—is essential to fully understand its capabilities and boundaries. A notable limitation is the increased training time compared to residual architectures. While this may contribute to stronger representations, optimizing training efficiency remains an open challenge, warranting further exploration of scheduling and initialization strategies.
By enabling implicit layer pruning, ACNs aim to reduce inference-time resource use, supporting more sustainable AI. Long-term, this could inspire adaptive architectures with System 1 / System 2 behavior (kahneman2011thinking, ), dynamically adjusting depth based on task complexity. However, like all efficiency advances, broader deployment may introduce ethical concerns—such as misuse in low-regulation environments or privacy-sensitive applications—highlighting the need for responsible development and oversight.
10 Acknowledgments and funding disclosure
This work has been partially supported by project MIS-5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the EU under the NextGenerationEU Program. We acknowledge EuroHPC JU project ID EHPC-AI-2024-A04-051 for use of the supercomputer LEONARDO@ CINECA, Italy.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main scope/contribution of the paper that is stated in the abstract and introduction is the introduction of an alternative to residual networks architecture, theoretically and empirically analyzed throughout the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a "Limitations" Section in the main paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In our theoretical results we provide assumptions (aka 1d linear analysis of gradient equations) and analytical derivations can be found in Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We clearly describe experiments and evaluation setup and we further provide all experimental details in the Appendix. Moreover, we provide an anonymous link where code for the paper will be released.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide an anonymous link where code for the paper will be released. Datasets used in the paper are publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In the main paper we discuss core details regarding experiments and in the Appendix we report detailed choices.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer:[Yes]
  34. Justification: We report average accuracy where applicable (e.g., noise experiments), and include mean metrics with standard deviation in continual learning settings where variability is significant.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Our experiments are small-scale but we consistently report training epochs.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer:[Yes]
  44. We conform to the NeurIPS Code of Ethics. We clearly describe our methods, include details to reproduce them and plan to release our code.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We include a "Broader Impact" section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer:[Yes]
  54. Justification: We do not release new data and we train small-scale discriminative models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We check the license and cite models and datasets we use in the experiments.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We train small reproducible models with public data and explain in detail our setups. We do not introduce major assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: not invloved.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: not involved
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs are only used for grammar and formatting.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Technical Appendices and Supplementary Material
Appendix A Extended Literature Review
We present here an extended literature review across four key areas relevant to our work: 1) residual connections and their role in training stability, 2) architectural variants with longer/denser residual connections, 3) methods that employ intermediate layer losses to learn better representations and exit early, and 4) neural architectures that induce regularization and better representations. Our work is more closely related to research area 4, but it is important to note that training stability, efficiency, performance, and representation learning are related goals, which can be achieved either through architectural choices (1, 2, 4) and / or through additional optimization criteria (3).
Residual Connections and Training Stability: Training deep neural networks with gradient descent becomes increasingly difficult as network depth increases. Multipath network architectures date back to the 1980s, with early work exploring cascade structures in fully connected networks trained layer by layer to improve training stability [15]. Highway networks [47] introduced gated bypass paths that allowed for effective training of networks with hundreds of layers. [20] found that deeper convolutional neural networks (CNNs) not only suffer from a decrease in generalization performance, often due to overfitting, but also experience a decrease in training performance. To address this, they introduced residual connections (or identity mappings), proposing that learning residual functions relative to identity mappings simplifies optimization. These skip connections improve the training process and often improve performance ([5], [36], [60], [32]). [53] further argued that a residual network with $n$ layers can be viewed as a collection of $2^{n}$ paths of varying length. At each layer, the signal either skips the layer or passes through it, creating $2^{n}$ possible paths. Despite sharing weights, these paths function as an ensemble of networks, as confirmed by experiments. In contrast, a traditional deep feedforward network has only one path, so removing any random layer significantly degrades performance. Additionally, the authors showed that these paths are typically shallow, with backward gradients often vanishing after passing through only a small fraction of the total layers.
Residual Variants: Following the success of residual networks various architectural modifications were proposed to improve efficiency and performance. In DenseNets [22], each layer is connected to all subsequent layers enabling for more efficient feature reuse; fusion is achieved through concatenation rather than addition. More recently, DenseFormer [37] introduced learned weighted averaging across layer outputs, while Depth-Wise Attention [14] applies attention mechanisms across block outputs.
Intermediate Supervision and Early Exit: In deeply supervised nets [31], complementary objectives are added to all intermediate layers to encourage hidden layers to learn more discriminative representations. In this approach, each intermediate objective $i$ is a loss function that captures the classification error of an SVM trained on the output features of layer $i$. The overall loss is the sum of the intermediate and final objectives. This idea evolved in several directions: Graves [17] proposed adaptive computation time, while more recent work like MSDNet [21] and CALM [44] introduced dedicated prediction heads. Other approaches employ trainable routing mechanisms [56, 57] to determine layer usage. Concurrent to our work, LayerSkip [12] proposes an architecture similar to ACNs, focusing primarily on inference acceleration through layer dropout and early exit mechanisms. Additionally, [24] also incorporates intermediate losses with a common head and a linearly increasing weight curriculum, justifying it through the lens of representational similarity between intermediate layers. While these approaches rely on explicit auxiliary objectives or dedicated components, our work achieves similar benefits through architectural design alone, enabling natural depth determination through gradient-based optimization.
Architecturally-induced regularization and representation learning: Stochastic regularization methods like Dropout [46] and its variants demonstrated that randomly dropping connections during training can lead to more robust feature learning. This insight was extended to other structural approaches like Stochastic Depth [23] where randomly dropping entire layers improved generalization. Residual connections initially proposed to address the vanishing gradient problem [20] have been shown to contribute to smoother loss landscapes and improved generalization [32]. These findings align with theoretical work showing that architectural choices impose implicit biases that influence the solutions found during training [19]. Recent work on transformers shows that architectural choices like attention patterns and layer normalization can also induce implicit regularization effects, e.g., the combination of skip connections and layer normalization can bias the model toward low-rank solutions [4]. The proposed long connection approach builds on these insights, using architectural design to naturally encourage the learning of robust representation while enabling automatic information compression allowing for early exit.
Appendix B Experimental Details and Setup
@@
CIFAR-10 - MLP Mixer: The MLP Mixers have 16 layers with a hidden size of 128. The patch size is 4 (the input is 32x32, 3 channels). The MLP dimension $D_{C}$ is $512$, while $D_{S}$ is $64$. We are using the AdamW optimizer [34] with a maximum learning rate of $0.001$ and a Cosine Scheduler with Warmup. The batch size is 64.
@@
Continual Learning Experiments: We are using the same MLP-Mixer setup with the CIFAR-10 experiment (see above). We train for 10 epochs in each task, using AdamW with learning rate of 0.001 and a batch size of 64. For Synaptic Intelligence we use a coefficient $\lambda=1$.
Transfer Learning experiment Hyperparameter Choices: In our main paper experiment we compare: 1) our proposed AC-Mixer, 2) a Residual Mixer with the setup of [24] (Aligned) and 3) a Residual Mixer with the setup of [13] (LayerSkip), with the rotational early exit curriculum with $p_{max}=0.1$, $e_{scale}=0.2$.
Appendix C Gradient Propagation equations derivation